OpenStudy (logan13):

Which is equal to ? A. 8−50 B. 8−15 C. 8−5 D. 8−2

OpenStudy (texaschic101):

x*m / x^n = x^m-n 8^-10 / 8^5 = 8^(-10 - 5) = 8^-15